# Understanding the Application of Utility Theory in Robotics and Artificial Intelligence: A Survey

### v1

@@ III Single-Agent Systems
For single-agent systems, cognitive robotics is the most typical research area for implementing Utility Theory. It studies the mechanisms, architectures, and constraints that allow lifelong and open-ended improvement of perceptual, reasoning, planning, social, knowledge acquisition, and decision-making skills in embodied machines [12]. Especially building a value  [^3] In this section, we use the terms value and utility interchangeably. system to mimic the “brain” of an AI agent mapping behavioral responses for sensed external phenomena is the core component of cognitive robotics, which is also an emerging and specialized sub-field in neurorobotics, robotics, and artificial cognitive systems research.
Here, the value measures an agent’s effort to expend to obtain a reward or avoid punishment. It is not hard-wired for an AI agent, even a biological entity, and the specific value system achieved through experience reflects an agent’s subjective evaluation of the sensory space. And the value mechanisms usually have been defined as the expected values, particularly in uncertain environments. Moreover, the innate value reflects an agent’s subjective evaluation of the sensory space, but the acquired value is shaped through experience during its development.
[25] defines the artificial value system for the autonomous robot as a mechanism that can reflect the effect of their experience in future behaviors. In the life of an agent, the agent starts from inherent values, and its acquired values are modified through interaction with various environments and knowledge accumulation in learning. So the acquired value is thus activity-dependent and allows the value system to become sensitive to stimuli that are not able to trigger a value-related response by themselves [26].
@@
According to the development of cognitive robotics, the existing value systems can mainly be classified into three categories. Neuroanatomical Systems discuss the explainable biologically inspired value systems design from neuroanatomy and physiology perspectives [27]; Neural Networks Systems build more abstract models through mathematical approaches to mimic the agent’s value systems; Motivational Systems consider the model that agents interact with environments to satisfy their innate values, and the typical mechanism is reinforcement learning (RL). We review the three categories as follow:
Figure 4: The illustration of the functions of the brain.
III-A Neuroanatomical Systems
@@
Especially based on the principles of neuromodulation in the mammalian brain, [31] presents a strategy for controlling autonomous robots. They used a cognitive robot – CARL-1, to test the hypothesis that neuromodulatory activity can shape learning, drive attention, and select actions. The experiments demonstrated that the robot could learn to approach and move away from stimuli by predicting the positive or negative value.
Recently, [32] shows the putative functions ascribed to dopamine (DA) emerging from the combination of a standard computational mechanism coupled with differential sensitivity to the presence of DA across the striatum through testing on the simulated humanoid robot iCub interacting with a mechatronic board. However, it is worth mentioning that the current trend tries to blur the line between designing value systems that follow the neuroanatomy and biology of the brain and artificial neural networks (ANN) that are biologically inspired but less anatomically accurate [12].
@@ III-B Artificial Neural Networks Systems
Especially [39] proposes a neural network architecture for integrating different value systems with reinforcement learning, presenting an empirical evaluation and comparison of four value systems for motivating exploration by a Lego Mindstorms NXT robot. [40] introduces a model of hierarchical curiosity loops for an autonomous active learning agent by selecting the optimal action that maximizes the agent’s learning of sensory-motor correlations. From the information theory perspective, [41, 42] presents a method linking information theoretic quantities on the behavioral level (sensor values) to explicit dynamical rules on the internal level (synaptic weights) in a systematic way. They studied an intrinsic motivation system for behavioral self-exploration based on the maximization of the predictive information using a range of real robots, such as the humanoid and Stumpy robots.
@@
[43] recently proposed the value system of a developmental robot using the RBM as the data structure. They use simulation to demonstrate the mechanism allowing the agent to accumulate knowledge in an organized sequence with gradually increasing complexity while hardly learning from purely random areas. From another angle, RL gradually becomes a popular and effective model for non-specific reward tasks [12], especially combined with deep learning methods.
III-C Motivational Systems
@@
To clarify agents’ motivation, providing an expected utility for each perceptual state in the motivation evaluation is the pre-condition in their decision-making, especially in discovering goals achieving them automatically, and determining the priority of drives [11]. Many machine learning algorithms share the properties of value systems and provide a way for AI agents to improve their performance in tasks and respect their innate values [12]. The most typical and general model is reinforcement learning (RL).
@@
The essence of RL is learning from interaction. When an RL agent interacts with the environment, it can observe the consequence of its actions and learn to change its behaviors based on the corresponding rewards received. Moreover, the theoretical foundation of RL is the paradigm of trial-and-error learning rooted in behaviorist psychology [23]. In this process, the best action sequences of the agent is determined by the rewards through interaction and the goal of the agent is to learn a policy $\pi$ maximizing its expected return (cumulative utilities with discount). In a word, given a state, a policy returns an action performance and learn an optimal policy from the consequence of actions through trial and error to maximize the expected return in the environment. Figure 1 illustrates this perception-action-learning loop.
Figure 6: The illustration of Agent Needs Hierarchy as five different modules.
@@
Furthermore, the flexibility of RL permits it as an effective utility system to model agents’ motivation through designing task-nonspecific reward signals derived from agents’ experiences [39]. [44] introduced one of the earliest self-motivated value systems integrated with RL, demonstrating that the SAIL robot could learn to pay attention to salient visual stimuli while neglecting unimportant input. To understand the origins of reward and affective systems, [45] built artificial agents sharing the same intrinsic constraints as natural agents: self-preservation and self-reproduction. [46] uses a table-based Q-learning as a baseline compared with other function approximation techniques in motivated reinforcement learning.
@@
From the mechanism of intelligent adaptive curiosity perspective, [47] introduced an intrinsic motivation system pushing a robot toward situations to maximize its learning progress. [48] presented a curiosity-driven reinforcement learning approach that explores the iCub’s state-action space through information gain maximization, learning a world model from experience, and controlling the actual iCub hardware in real time. To help the iCub robot intrinsically motivated to acquire, store and reuse skills, [49] introduced Continual Curiosity driven Skill Acquisition (CCSA), a set of compact low-dimensional representations of the streams of high-dimensional visual information learned through incremental slow feature analysis.
@@
Especially [50] proposed an RL model combined with the agent basic needs satisfaction and intrinsic motivations. It was tested in a simulated survival domain, where a robot was engaged in survival tasks such as finding food or water while avoiding dangerous situations. Furthermore, [51, 52] recently proposed the first formalized utility paradigm – Agent Needs Hierarchy – for the needs of an AI agent to describe its motivations from the system theory and psychological perspective. The Agent Needs Hierarchy is similar to Maslow’s human needs pyramid [15]. An agent’s abstract needs for a given task are prioritized and distributed into multiple levels, each of them preconditioned on their lower levels. At each level, it expresses the needs as an expectation over the corresponding factors/features’ distribution to the specific group [52]. Fig. 6 illustrates the Agent Needs Hierarchy as five different modules in the proposed Self-Adaptive Swarm Systems (SASS) [53].
Specifically, the lowest (first) level is the safety features of the agent (e.g., features such as collision detection, fault detection, etc., that assure safety to the agent, human, and other friendly agents in the environment). The safety needs can be calculated through its safety feature’s value and corresponding safety feature’s probability based on the current state of the agent. After satisfying safety needs, the agent considers its basic needs, which includes features such as energy levels, data communication levels that help maintain the basic operations of that agent. Only after fitting the safety and basic needs, an agent can consider its capability needs, which are composed of features such as its health level, computing (e.g., storage, performance), physical functionalities (e.g., resources, manipulation), etc.
At the next higher level, the agent can identify its teaming needs that accounts the contributions of this agent to its team through several factors (e.g., heterogeneity, trust, actions) that the team needs so that they can form a reliable and robust team to successfully perform a given mission.
Ultimately, at the highest level, the agent learns skills or features to improve its capabilities and performance in achieving a given task through RL. Later work by [54] presented a strategy-oriented Bayesian soft actor-critic (BSAC) Model based on the agent needs, which integrates an agent strategy composition approach termed Bayesian Strategy Network (BSN) with the soft actor-critic (SAC) method to achieve efficient deep reinforcement learning (DRL).
Generally, the set of agent needs reflects its innate values and motivations. It is described as the union of needs at all the levels in the needs hierarchy. Each category needs level is combined with various similar needs (expected utilities) presenting as a set, consisting of individual hierarchical and compound needs matrix [55]. It drives the agent to develop various strategies and diverse skills to satisfy its intrinsic needs and values, much like a human does.
IV Multi-Agent Systems
@@
However, in the MAS setting, the situations will become much more complex when we consider individual utility reflects its own needs and preferences. For example, although we assume each group member receives the same team rewards in fully cooperative MAS, the benefits received by an individual agent are usually significant differences according to its contributions and innate values in real-world scenarios or general multi-agent settings. Especially in distributed intelligent systems, such as multi-robot systems (MRS), self-driving cars, and delivery drones, there is no single decision-maker with full information and authority. Instead, the system performance greatly depends on the decisions made by interacting entities with local information and limited communication capabilities [60]. This section reviews the works of decision-makers compromise in MAS from the game-theoretic utility cooperative decision-making perspective based on their needs and innate values. Then, we discuss the design and influence of individual utility functions to induce desirable system behaviors and optimization criteria.
IV-A Game-theoretic Utility Systems
From a game-theoretic control perspective, most of the research in game theory has been focused on single-stage games with fixed, known agent utilities [61], such as distributed control in communication [62] and task allocation [63]. Especially, recent MAS research domains focus on solving path planning problems for avoiding static or dynamical obstacles [64] and formation control [65] from the unintentional adversary perspective. For intentional adversaries, the Pursuit Domain [66, 67] primarily deals with how to guide pursuers to catch evaders [68] efficiently. Similarly, the robot soccer domain [69] deals with how one group of robots wins over another group of robots on a typical game element.
@@ IV-A Game-theoretic Utility Systems
More recently, [80, 81] proposed a new model called the Game-Theoretic Utility Tree (GUT), combining the core principles of game theory and utility theory to achieve cooperative decision-making for MAS in adversarial environments. It combines with a new payoff measure based on agent needs for real-time strategy games. By calculating the Game-Theoretic Utility Computation Unit distributed at each level, the individual can decompose high-level strategies into executable low-level actions. It demonstrated the GUT in the proposed Explore Game domain and verified the effectiveness of the GUT in the real robot testbed – Robotarium [82], indicating that GUT can organize more complex relationships among MAS cooperation and help the group achieve challenging tasks with lower costs and higher winning rates. Fig 7 illustrates a MAS (UAV) working in adversarial environments with several typical scenarios.
IV-B Cooperative Decision-making and Learning
Figure 7: Illustration of A MAS (UAV) Working in Adversarial Environments with Several Typical Scenarios

### v3

@@ 3 Single-Agent Systems
For single-agent systems, cognitive robotics is the most typical research area for implementing Utility Theory. It studies the mechanisms, architectures, and constraints that allow lifelong and open-ended improvement of perceptual, reasoning, planning, social, knowledge acquisition, and decision-making skills in embodied machines [173]. Especially building a value system to mimic the “brain” of an AI agent mapping behavioral responses for sensed external phenomena is the core component of cognitive robotics, which is also an emerging and specialized sub-field in neurorobotics, robotics, and artificial cognitive systems research.
@@
From a neuroeconomic perspective, a variety of cortical and subcortical areas are associated with neural firing of reward-predictive and reward-prediction error signals, including the orbitofrontal cortex, prefrontal cortex, parietal cortex, and striatum. Fig. 13 illustrates a schematic model of reinforcement learning implementation in the cortico-basal ganglia circuit [96] and the various aspects of value system development from a biological brain to an artificial brain, including external observation, internal structure, infrastructure, and social structure [151]. Similarly, to build a utility-based reward system in robots, existing value systems can be classified into three categories: Neuroanatomical Systems discuss the explainable biologically inspired value systems design from neuroanatomy and physiology perspectives [238]; Neural Networks Systems build more abstract models through mathematical approaches to mimic the agent’s value systems; Motivational Systems consider the model that agents interact with environments to satisfy their innate values, and the typical mechanism is reinforcement learning (RL). We review the three categories as follow:
3.1 Neuroanatomical Systems
@@
Figure 14: The illustration of the functions of the brain.
Especially based on the principles of neuromodulation in the mammalian brain, [63] presents a strategy for controlling autonomous robots. They used a cognitive robot – CARL-1, to test the hypothesis that neuromodulatory activity can shape learning, drive attention, and select actions. The experiments demonstrated that the robot could learn to approach and move away from stimuli by predicting the positive or negative value. Moreover, [80] shows the putative functions ascribed to dopamine (DA) emerging from the combination of a standard computational mechanism coupled with differential sensitivity to the presence of DA across the striatum through testing on the simulated humanoid robot iCub interacting with a mechatronic board.
$\begin{split}\frac{dx}{dt}=-Lx+I\end{split}$ (10)
To summarize, although researchers have proposed numerous neuroanatomical models, the most common node type is the leaky integrator neural unit. Leaky integrator units have the differential equations Eq. (10). Where $x$ is the activation potential of the neuron, $I$ is the input to the neuron, $L$ is the rate of “leak” of potential, and $t$ is time. Additionally, researchers use leaky integrator units to model intrinsic motivations, such as novelty, in cognitive models of value systems. It is worth mentioning that the current trend tries to blur the line between designing value systems that follow the neuroanatomy and biology of the brain and artificial neural networks (ANN) that are biologically inspired but less anatomically accurate [173]. The following section examines the ANN as a value system.
3.2 Artificial Neural Networks Value Systems
@@
Especially [174] proposes a neural network architecture for integrating different value systems with reinforcement learning, presenting an empirical evaluation and comparison of four value systems for motivating exploration by a Lego Mindstorms NXT robot. The value system is a multi-layer neural network. The first layers process sensations to transmit information to motivation layers for computing motivation values, and the reinforcement learning layers compute weight values to generate actions and update weights through Q-learning. The framework has a combined memory model for the value system and RL components, as well as computational motivation models to guide exploration and learning. Fig. 16 shows the generic neural network structure used to compare different motivations in a value system. Also, [102] introduces a model of hierarchical curiosity loops for an autonomous active learning agent by selecting the optimal action that maximizes the agent’s learning of sensory-motor correlations.
@@
From the information theory perspective, [161, 162] presents a method linking information theoretic quantities on the behavioral level (sensor values) to explicit dynamical rules on the internal level (synaptic weights) in a systematic way. They studied an intrinsic motivation system for behavioral self-exploration based on the maximization of the predictive information using a range of real robots, such as the humanoid and Stumpy robots. Additionally, [156] proposed the value system of a developmental robot using the RBM as the data structure. They use simulation to demonstrate the mechanism allowing the agent to accumulate knowledge in an organized sequence with gradually increasing complexity while hardly learning from purely random areas. As we can see, RL has become a popular and effective model for non-specific reward tasks, especially when combined with deep learning methods implemented in motivational systems modeling, which we discuss in the following part.
Figure 16: The illustration of the generic neural network structure used to compare different motivations in a value system [174].
3.3 Motivational Systems
@@
To clarify agents’ motivation, providing an expected utility for each perceptual state in the motivation evaluation is the pre-condition in their decision-making, especially in discovering goals achieving them automatically, and determining the priority of drives [220]. Many machine learning algorithms share the properties of value systems and provide a way for AI agents to improve their performance in tasks and respect their innate values [173].
3.3.1 Reinforcement Learning
Figure 17: Rewards (Cumulative Utilities) Driven Behaviors
@@ 3.3.1 Reinforcement Learning
The most typical and general model is reinforcement learning (RL). It is an excellent model to describe the needs-driven behaviors of AI agents. The essence of RL is learning from interaction based on reward-driven (such as utilities and needs) behaviors, much like natural agents. When an RL agent interacts with the environment, it can observe the consequence of its actions and learn to change its behaviors based on the corresponding rewards received. The theoretical foundation of RL is the paradigm of trial-and-error learning rooted in behaviorist psychology [246]. In RL, the policy dictates the actions that the agent takes as a function of its state and the environment, and the goal of the agent is to learn a policy maximizing the expected cumulative rewards/utilities in the process. Fig. 17 illustrates the RL process of rewards-driven behaviors [262].
When an RL agent interacts with the environment, it can observe the consequence of its actions and learn to change its behaviors based on the corresponding rewards received. Moreover, the theoretical foundation of RL is the paradigm of trial-and-error learning rooted in behaviorist psychology [246]. In this process, the best action sequences of the agent is determined by the rewards through interaction and the goal of the agent is to learn a policy $\pi$ maximizing its expected return (cumulative utilities with discount). In a word, given a state, a policy returns an action performance and learn an optimal policy from the consequence of actions through trial and error to maximize the expected return in the environment. Figure 17 illustrates this perception-action-learning loop.
3.3.2 Integrating a Robot Utility Value System to RL
@@
The flexibility of RL permits it as an effective utility system to model agents’ motivation through designing task-nonspecific reward signals derived from agents’ experiences [174]. [111] introduced one of the earliest self-motivated value systems integrated with RL, demonstrating that the SAIL robot could learn to pay attention to salient visual stimuli while neglecting unimportant input. To understand the origins of reward and affective systems, [71] built artificial agents sharing the same intrinsic constraints as natural agents: self-preservation and self-reproduction. [175] uses a table-based Q-learning as a baseline compared with other function approximation techniques in motivated reinforcement learning.
@@
From the mechanism of intelligent adaptive curiosity perspective, [195] introduced an intrinsic motivation system pushing a robot toward situations to maximize its learning progress. [85] presented a curiosity-driven reinforcement learning approach that explores the iCub’s state-action space through information gain maximization, learning a world model from experience, and controlling the actual iCub hardware in real-time. To help the iCub robot intrinsically motivated to acquire, store and reuse skills, [124] introduced Continual Curiosity driven Skill Acquisition (CCSA), a set of compact low-dimensional representations of the streams of high-dimensional visual information learned through incremental slow feature analysis. RL agents learn various skills by associating intrinsic rewards with improvements to the world model [150]. The acquired skill includes both the learned actions and the feature representation, which greatly change the feature output. Moreover, these skills will be stored and reused to generate new observations for continually developing advanced skills.
Figure 18: The illustration of Robot Needs Hierarchy as five different levels [282].
@@
Additionally, [69] proposed an RL model combined with the agent basic needs satisfaction and intrinsic motivations. It was tested in a simulated survival domain, where a robot was engaged in survival tasks such as finding food or water while avoiding dangerous situations. Furthermore, [282, 283] recently proposed the first formalized utility paradigm – Agent Needs Hierarchy – for the needs of an AI agent to describe its motivations from the system theory and psychological perspective. The Agent Needs Hierarchy is similar to Maslow’s human needs pyramid [163]. An agent’s abstract needs for a given task are prioritized and distributed into multiple levels, each of them preconditioned on their lower levels. At each level, it expresses the needs as an expectation over the corresponding factors/features’ distribution to the specific group [283]. Ultimately, at the highest level, the agent learns skills or features to improve its capabilities and performance in achieving a given task through RL. Fig. 18 illustrates the Agent Needs Hierarchy as five different levels in the proposed Self-Adaptive Swarm Systems (SASS).
Later work by [287, 288] presented a strategy-oriented Bayesian soft actor-critic (BSAC) Model based on the agent needs, which integrates an agent strategy composition approach termed Bayesian Strategy Network (BSN) with the soft actor-critic (SAC) method to achieve efficient deep reinforcement learning (DRL). The key idea of the BSAC is that by building a suitable BSN to describe complicated relationships among subsystems in the complex system, it can transform a complex policy into a joint policy consisting of several simple sub-policies, which can improve the sample efficiency and training speed. From another perspective, the subsystems’ policy training is like muscle memory training in human skills development. Fig. 19 illustrates a Bayesian Strategy Network for a robot dog model that implements the Actor-Critic DRL architecture.
Furthremore, [279] propose a new RL model termed innate-values-driven RL (IVRL) based on combined motivations’ models and expected utility theory to mimic its complex behaviors in the evolution through decision-making and learning. Compared with the traditional RL model, IVRL generates the input state and rewards from the critic module instead of directly from the environment, which means that the AI agent receives various utilities from the environment through executing an action or strategy in the IVRL model. Moreover, the individual needs to calculate innate values (expected utility) through its needs weights and current utilities and then select suitable actions or strategies to optimize or maximize its accumulated expected utility. Fig. 20 illustrates the IVRL model. By adjusting needs weights in its innate-values system, it can adapt to different tasks representing corresponding characteristics to maximize the rewards efficiently (Eq. (11)). In other words, through interacting with other agents and environments, the agent builds its unique value system to adapt to them, the same as "Individual is the product of their own environment." Generally, the set of agent needs reflects its innate values and motivations. It is described as the union of needs at all levels in the needs hierarchy, forming the individual hierarchical and compound needs matrix, which drives the agent to develop various strategies and diverse skills to satisfy its intrinsic needs and values, much like a human does [277].
$\begin{split}Q^{*}(s,w,a)&=\max_{\omega,\pi}\mathbb{E}[G_{t}|S_{t}=s_{t},W_{t}=w_{t},A_{t}=a_{t},\omega,\pi]\\ &=\mathbb{E}_{s^{\prime}\sim\epsilon}\left[r+\gamma\max_{w^{\prime},a^{\prime}}Q^{*}(s^{\prime},w^{\prime},a^{\prime})\bigg|s,w,a\right]\end{split}$ (11)
Figure 19: An Bayesian Strategy Network model of a robot dog implementing the Actor-Critic DRL architecture model [287].
To summarize, when integrating value systems into RL, we consider model-based and model-free learning alongside exemplar and function approximation learning. Two commonly used value updates modify either a utility value associated with the expected future reward of being in a given state (Eq. (12)) or a utility value associated with the expected future reward of taking a given action in a given state (Eq. (13)) [173]. Where the value updates involve multiple utility reward types, including intrinsic (novelty, curiosity, creativity, and competence/learning progress) and extrinsic (fear, thirst, and hunger).
$\begin{split}V_{t}(S_{t-1})=&V_{t-1}(S_{t-1})+\\ &\alpha\left[R_{t}+\gamma V_{t-1}(S_{t})-V_{t-1}(S_{t-1})\right];\end{split}$ (12)
$\begin{split}&Q_{t}(S_{t-1}-A_{t-1})=Q_{t-1}(S_{t-1}-A_{t-1})+\\ &\alpha\left[R_{t}+\gamma\max_{A\in\mathcal{A}}Q_{t-1}(S_{t}-A)-Q_{t-1}(S_{t-1}-A_{t-1})\right]\end{split}$ (13)
4 Multi-Agent Systems
@@
However, in the MAS setting, the situations will become much more complex when we consider individual utility reflects its own needs and preferences. For example, although we assume each group member receives the same team rewards in fully cooperative MAS, the benefits received by an individual agent are usually significant differences according to its contributions and innate values in real-world scenarios or general multi-agent settings. Especially in distributed intelligent systems, such as multi-robot systems (MRS), self-driving cars, and delivery drones, there is no single decision-maker with full information and authority. Instead, the system performance greatly depends on the decisions made by interacting entities with local information and limited communication capabilities [197].
From the multi-objective multi-agent perspective, [213] proposes a taxonomy based on utility into three types: Individual utility, where each agent serves a different agenda and just tries to optimise for that; Team utility, where all agents serve the same interest, like playing football; Social choice utility, which represents agents are interested in optimizing the overall social welfare across all agents. According to different application scenarios, they provided a more detailed classification. Fig. 21 illustrates the taxonomy of multi-objective multi-agent decision-making settings based on utility theory and Fig. 22 presents the utility rewards classification in multi-objective multi-agent decision-making with different setting.
This section, in combination with the applications in AI and robotics, reviews work on decision-makers’ compromise in MAS, ranging from game-theoretic utilities to cooperative decision-making grounded in their needs and innate values. Then, we discuss the design and the influence of individual utility functions on desirable system behaviors and optimization criteria. Finally, we extend the topic from individual utility to applications of artificial social welfare.
4.1 Game-theoretic Utility Systems
Figure 20: The illustration of the architecture of the IVRL Actor-Critic model [279].
In multi-objective decision-making, each agent aims to optimise its reward utility across a set of objectives and to make compromises between competing objectives based on its utility function. However, when multiple self-interested agents learn and act together in the same environment, it is generally not possible for all agents to receive the maximum possible reward. Therefore, MAS are often designed to converge to a Nash equilibrium [67].
From a game-theoretic control perspective, most of the research in game theory has been focused on single-stage games with fixed, known agent utilities [48], such as distributed control in communication [294] and task allocation [17]. Especially, recent MAS research domains focus on solving path planning problems for avoiding static or dynamical obstacles [3] and formation control [232] from the unintentional adversary perspective. For intentional adversaries, the Pursuit Domain [60, 125] primarily deals with how to guide pursuers to catch evaders [155] efficiently. Similarly, the robot soccer domain [183] deals with how one group of robots wins over another group of robots on a typical game element. Fig 23 illustrates a group of Unmanned Aerial Vehicles (UAV) working in adversarial environments with several typical scenarios.
@@ 4.1 Game-theoretic Utility Systems
More recently, [286, 285] proposed a new model called the Game-Theoretic Utility Tree (GUT), combining the core principles of game theory and utility theory to achieve cooperative decision-making for MAS in adversarial environments. It combines with a new payoff measure based on agent needs for real-time strategy games. By calculating the Game-Theoretic Utility Computation Unit distributed at each level, the individual can decompose high-level strategies into executable low-level actions. It demonstrated the GUT in the proposed Explore Game domain and verified the effectiveness of the GUT in the real robot testbed – Robotarium [205], indicating that GUT can organize more complex relationships among MAS cooperation and help the group achieve challenging tasks with lower costs and higher winning rates. Fig. 24 demonstrates five pursuers catching one evader in the Pursuit-Evasion game through the game-theoretic utility tree (GUT) with circle formation (level 1) and PP tactic (level 2) in Robotarium.
4.2 Cooperative Decision-making and Learning
Figure 21: The illustration of the taxonomy of multi-objective multi-agent decision making settings [213].